OpenStudy (anonymous):

Can someone show me how to solve this? Thank you :) How many solution does system? -3x+6y=10 -3x+6y=-4

OpenStudy (campbell_st):

well they are the same values for x and y on the left... that equal different values on the right... it doesn't work...

OpenStudy (anonymous):

I know that's why i'm confused

OpenStudy (anonymous):

if it helps i'm doing a study guide for solving system of equation u elimination

OpenStudy (anonymous):

and substitution

OpenStudy (anonymous):

by*

OpenStudy (anonymous):

its not possible because if you multiply both sides of one of the equations by -1 you get: -1(-3x+6y)=-1(10) which equals 3x-6y=-10 so then you add them together (i cant remember the name of this method...) to be:  and this is not possible because 0 does not equal 14